According to a​ study, 47​% of all males between the ages of 18 and 24 live at home. ​ (Unmarried college students living in a dorm are counted as living at​ home.) Suppose that a survey is administered and 125 of 234 respondents indicated that they live at home.​ (a) Use the normal approximation to the binomial to approximate the probability that at least 125 respondents live at home.​ (b) Do the results from part​ (a) contradict the​ study?
Answer #1

Answer: According to a​ study, 47​% of all males between the ages of 18 and 24 live at home. ​ (Unmarried college students living in a dorm are counted as living at​ home.) Suppose that a survey is administered and 125 of 234 respondents indicated that they live at home.​

Solution:

p = 47% = 0.47

n = 234

X = 125

(a) Use the normal approximation to the binomial to approximate the probability that at least 125 respondents live at home.​

X ~ Binomial (μ, σ)

μ = np = 234 * 0.47 = 109.98

σ = √np(1-p) = √234*0.47*(1-0.47) = 7.6348

P(X ≥ 125) = P(X ≥ 124.5)

P(X ≥ 125) = P(Z ≥ 124.5 - 109.98 / 7.6348)

P(X ≥ 125) = P(Z ≥ 1.9018)

P(X ≥ 125) = 1 - P(Z < 1.9018)

P(X ≥ 125) = 1 - 0.9714

P(X ≥ 125) = 0.0286

(b) Do the results from part​ (a) contradict the​ study?

Yes, because the probability in part (a) is less than 0.05.